OpenStudy (anonymous):

Which function is an odd function? f(x) = 5x2 f(x) = 3x – 1 f(x) = 0.7x2 + 3 f(x) = 0.8x3

OpenStudy (marcelie):

last one ?

OpenStudy (astrophysics):

It is an even function if f(-x) = f(x), it is an odd function if f(-x) = - f(x)

OpenStudy (astrophysics):

That's one way to test it!

OpenStudy (marcelie):

isnt odd functions mostly cubic?

OpenStudy (astrophysics):

Correct

OpenStudy (marcelie):

im guessing its the last one

OpenStudy (astrophysics):

You can think of this for example, say we're given \[(-2)^3 = (-2)(-2)(-2) = -8\]

OpenStudy (astrophysics):

Notice how it's still negative

OpenStudy (astrophysics):

Hope that helps!